Triple Intergrals

By using triple integrals, find the volume of the solid G bounded above by plane 2x+y+z=6 and below the plane z=1 in the first octant
Answer #1

plam. : in first octent V=SSSav =SSS dz dy dn (1-- 9) dy dn 1-- %3D -n+2x dn [x-x*+), %3D dbm 우 =(우+1-1)후un lz- Sf= xp fip (A-D %3D %3D %3D fenits go sLy plane, y2-pdon O Y+2 =1, y =x?, 9+2=1, y=x? V= JSS dv = SSU-3)AR. S S(I-3)dR2= 4. y=9 2=0, R. when R 4 dR dx dy = 4 2 3 dy 3/2 16 %3D *27 = 194 %3D %3D V= 194 (enit}3 2x+4+2 =9, 2=0, y-0, ssca.ze-3)48.

answered by: ANURANJAN SARSAM